OpenStudy (anonymous):

d/dx (3x^4-3x)/(3x^4+3x)

OpenStudy (turingtest):

the quotient rule is d/dx(u/v)=(u'v-v'u)/v^2 here we have u=(3x^4-3x) v=(3x^4+3x) find u' and v' and plug them into the formula

OpenStudy (anonymous):

12x^3-3 = u' and 12x^3+3 = v"

OpenStudy (turingtest):

perfect, but it's just v' not v'' now plug into the formula for the product rule and simplify

OpenStudy (anonymous):

i meant v' lol